# Ill-posedness of the Dirichlet problem for 2D Lagrangian mean curvature equation

### v1

@@ Abstract.
We investigate the Dirichlet problem of the two dimensional Lagrangian mean curvature equation in a bounded domain. Infinitely many $C^{1,\alpha}(\alpha\in(0,\frac{1}{5}))$ very weak solutions are built through Nash-Kuiper construction. Moreover, we note there are infinitely many $C^{1,\alpha}$ very weak solutions that can not be improved to be $C^{2,\alpha}$.
@@ 1. Introduction
In the present paper, we study the Dirichlet problem for two dimensional Lagrangian mean curvature equation
@@
where $\Omega\subset\operatorname{\mathbb{R}}^{2}$ is a bounded domain, phase function $\Theta:\overline{\Omega}\to(-\pi,\pi)$ is a given $C^{2,\kappa}(\kappa\in(0,1))$ function. The equation
@@
(1.3) $\arctan\lambda_{1}+\arctan\lambda_{2}=\Theta,$
@@
Generally, $n$-dimensional Lagrangian mean curvature equation is
@@
where $\lambda_{i}$’s are eigenvalues of $\nabla^{2}v$. The mean curvature vector of $n$-dimensional gradient graph $\Sigma$ is equal to $J\nabla_{g}\Theta$, where $g=\mathrm{Id}+(\nabla^{2}v)^{2}$ is the induced metric on $\Sigma$, and $J(x,y)=(-y,x)$ is a rotation in $\operatorname{\mathbb{R}}^{n}\times\operatorname{\mathbb{R}}^{n}$. In particular, when $\Theta$ is a constant, (1.4) is called the special Lagrangian equation, which originates in the special Lagrangian geometry introduced by Harvey-Lawson [14]. Note $\Sigma$ is a minimal surface in $\operatorname{\mathbb{R}}^{n}\times\operatorname{\mathbb{R}}^{n}$ if and only if $\Theta$ is a constant.
@@ 1.1. Background
In this paper, we consider not only classical solutions but also very weak solutions to (1.1). In fact, very weak solution to the Monge-Ampère equation was first introduced by Lewicka-Pakzad [18] via very weak Hessian defined by Iwaniec [16]. Two dimensional very weak Hessian coincides with the Monge-Ampère measure for any convex solution ([11]). Moreover, Without boundary conditions, Lewicka-Pakzad [18], Székelyhidi and the first author [7], and Hirsch-Inauen and the first author [5] proved that there are infinitely many very weak solutions to the two dimensional Monge-Ampère equation of different regularity. After imposing boundary data, the first author in [4] studied very weak solutions to the Monge-Ampère equation. Recently, Li-Qiu [19] also constructed infinitely many very weak solutions of the Dirichlet problem for 2-Hessian equation. All the above approaches to construct very weak solutions rely on the Nash-Kuiper construction, which is originated from isometric embeddings (Nash-Kuiper theorem, see [22, 17, 9, 6, 10, 23]), and developed to be convex integration to prove $h$-principle (see [13]). In recent decades, the technique of convex integration become a powerful approach to show non-uniqueness of nonlinear partial differential equations.
@@ 1.2. Main result and strategies
Along the line of very weak solutions to the Monge-Ampère equation, we also introduce the concept of very weak solution to the Dirichlet problem (1.1).
@@
$v\in C^{1}(\Omega)$ is a very weak solution to the Dirichlet problem (1.1) in $\Omega$, if $v=g$ on $\partial\Omega$, and it solves the following equation in the very weak sense in $\Omega$,
(1.5) $\cos\Theta\mbox{ curl curl }(v\cdot\mbox{Id})+\sin\Theta\left(-\frac{1}{2}\mbox{ curl curl }(\nabla v\otimes\nabla v)-1\right)=0,$
namely, for any $\phi\in C^{\infty}_{0}(\Omega)$
@@
[…]
@@
Let $\Omega\subset\operatorname{\mathbb{R}}^{2}$ be a bounded simply connected domain with $C^{2,\kappa}$ boundary $\partial\Omega$ for any given $\kappa\in(0,1)$. Suppose $g\in C^{2,\kappa}(\partial\Omega)$ and the phase $\Theta:\overline{\Omega}\to(-\pi,\pi)$ is a $C^{2,\kappa}$ function. Then the following statements hold.
@@
    If $|\sin\Theta|\leq c_{1}$ for some small positive constant $c_{1}$, then the Dirichlet problem (1.1) admits a unique $C^{2,\kappa}(\Omega)$ classical solution.
@@
    If $|\sin\Theta|\geq c_{2}$ for any positive constant $c_{2}\in(0,1]$, then the Dirichlet problem (1.1) admits infinitely many $C^{1,\alpha}(\Omega)$ very weak solutions for any $\alpha\in(0,\frac{1}{5})$.
@@
The assumption of simply connected is just from the application of Lemma A.2 in the proof of Proposition 3.1.
Note $c_{2}$ is arbitrary, as a quick observation, we have
@@
Our strategy to show Theorem 1.1 (1) is considering (1.2) as a perturbation of Laplacian equation and then applying the implicit function theorem. As for Theorem 1.1 (2), with $|\sin\Theta|\geq c_{2}>0$, the problem is reduced to solve
@@
We will solve the above equation by the Nash-Kuiper construction. The main obstacle is how to treat terms with phase function, comparing with [4], which will be regarded as a Laplacian of some function, namely,
$\cot\Theta\Delta v=\Delta(v\cot\Theta-V),\text{ in the very weak sense,}$,
@@
curlcurl( sym $\mathrm{curl}\,\mathrm{curl}\,\big{(}\mbox{ sym }\nabla w+\frac{1}{2}\nabla v\otimes\nabla v-(v\cot\Theta)\cdot\mathrm{Id}+V\cdot\mathrm{Id}\big{)}=-1,$
in the very weak sense, and then $V\cdot\textrm{Id}$ becomes an error term of lower order. With the elliptic estimate in Lemma 3.4, we are able to control it and arrive at our conclusion.
We remark that the method proving Theorem 1.1 (2) has more general application.
@@
where $a=\cot\tau$, $b=\sqrt{|\cot^{2}\tau-1|}$. With slight modifications, we can establish result like Theorem 1.1 (2) to the following Dirichlet problem
@@
where $\lambda_{1}$, $\lambda_{2}$ are eigenvalues of $\nabla^{2}v$. We note the equations $F_{\tau}(\lambda)=\Theta$ can be formulated to equations of the following form
@@
Equations in (1.7) are originated from the study of gradient graphs $\Sigma$. By Warren [26], if $\Sigma$ is a minimal Lagrangian graph in $(\operatorname{\mathbb{R}}^{2}\times\operatorname{\mathbb{R}}^{2},g_{\tau})$ with
$g_{\tau}=\sin\tau\big{(}\sum_{i=1}^{2}\mathrm{d}x_{i}\otimes\mathrm{d}x_{i}+\sum_{j=1}^{2}\mathrm{d}y_{j}\otimes\mathrm{d}y_{j}\big{)}+\cos\tau\big{(}\sum_{i=1}^{2}\mathrm{d}x_{i}\otimes\mathrm{d}y_{i}+\sum_{j=1}^{2}\mathrm{d}y_{j}\otimes\mathrm{d}x_{j}\big{)},$
@@
$F_{\tau}(\lambda(\nabla^{2}v))=\text{constant},$
We refer readers to results about gradient graphs in general dimensions in [26].
@@ 1.3. Organization and notations
$[h]_{m+\alpha}=\max_{|\beta|=m}\sup_{x\neq y}\frac{|\partial^{\beta}h(x)-\partial^{\beta}h(y)|}{|x-y|^{\alpha}}.$
Here $m$ is a nonnegative integer, $\alpha\in(0,1]$ a constant, and $\beta$ be a multi-index. We denote by $B_{r}$ the ball in $\operatorname{\mathbb{R}}^{2}$ with radius $r$ centering at 00. Let $\varphi\in C^{\infty}_{0}(B_{1})$ be a standard two dimensional mollifier. We define $\varphi_{l}(x)=l^{-2}\varphi(l^{-1}x)$, and $l>0$ is called the length-scale. We also denote
[…]
Some properties of mollification can be found in Lemma A.1. The operator $\mathrm{curl}$ in two dimension is defined by
[…]
Besides, for $\xi=(\xi_{1},\xi_{2})\in\mathbb{R}^{2}$ and $\eta=(\eta_{1},\eta_{2})\in\mathbb{R}^{2}$, we write $\xi\otimes\eta$ a $2\times 2$-matrix with entries $(\xi\otimes\eta)_{ij}=\xi_{i}\eta_{j}$. By a direct calculation, for a $C^{3}$-function $w$, we have
@@
Acknowledgement
Wentao Cao’s research is supported in part by the National Natural Science Foundation of China (Grant No. 12471224). Zhehui Wang’s research is supported in part by Young Scientists Fund of the National Natural Science Foundation of China (Grant No. 12401247), and Guangdong Basic and Applied Basic Research Foundation (Grant No. 2023A1515110910).
@@ Acknowledgement
The proof of Theorem 1.1 is based on an application of the implicit function theorem.
@@
Claim: There is a unique $\psi\in\mathcal{X}$ and a small enough constant $\mu\in(0,1)$ such that $G(\psi,\Psi)=0$ for any $\Psi$ satisfying $\|\Psi\|_{\kappa}\leq\mu$.
@@
for any $0<\sigma\leq r$ and $f\in C^{1,\kappa}(B_{r})$. Hence, if $|\sin\Theta|\leq c_{1}$ with small constant $c_{1}$ to be fixed, then we can get
@@
Then, $v:=\psi+\widetilde{g}$ uniquely solves
$&\Delta v+\tan\Theta(\det\nabla^{2}v-1)=0\mbox{ in }\Omega,\\ &v=g\mbox{ on }\partial\Omega.\end{aligned}\right.$
@@
3. Stage Proposition
The rest of the paper is devoted to show Theorem 1.1 (2). Similar to the treatment in [4, 7, 18] for the Monge-Ampère equation, we split (1.2) (after dividing $\sin\Theta$) into a differential system
(3.1) curl curl $-\mbox{curl curl }A=1,$
@@ 3. Stage Proposition
$\Delta V&=F(v),\mbox{ in }\Omega,\\ V&=0,\mbox{ on }\partial\Omega,\end{aligned}\right.$
@@
In fact, by a direct calculation, we note that the $v$ satisfying (3.1)-(3.2) solves (1.5) in the very weak sense.
To solve (3.2), our strategy is applying the technique of Nash-Kuiper construction to construct a sequence of subsolutions satisfying boundary condition, whose limit is our desired solution. Set
@@
Note that the definition is a slight modification of that in [4]. To construct new adapted subsolutions, we need to add the deficit $A-D(v,w)$ to the given adapted subsolution $v.$ Since the deficit is of the form $\rho^{2}(\mathrm{Id}+H)$, we require the following stage proposition (in the terminology of Nash [22]) for iteration, which is parallel to Proposition 3.1 in [4] and whose proof is different due to the new error terms. The new error terms are then controlled through Lemma 3.29.
@@
Suppose $\Sigma\subset\operatorname{\mathbb{R}}^{2}$ is a simply connected open bounded set with $C^{2}$ boundary. Let $v\in C^{2}(\overline{\Sigma})$ and $w\in C^{2}(\Sigma,\operatorname{\mathbb{R}}^{2})$ such that
@@
and let $\rho\in C^{1}(\overline{\Sigma})$, $H\in C^{1}(\overline{\Sigma},\operatorname{\mathbb{R}}^{2\times 2}_{sym})$ such that
@@
for some $\gamma>1$, $10<\delta,\,\alpha<$ 1 and $\lambda>1$ with $2\leq\lambda^{\alpha}$ and $\delta^{1/2}\lambda>1$. Then for any $\tau>1$, there exist functions $v^{\star}\in C^{2}(\overline{\Sigma})$, $w^{\star}\in C^{2}(\overline{\Sigma},\operatorname{\mathbb{R}}^{2})$ and symmetric matrix $\operatorname{\mathcal{E}}\in C^{1}(\overline{\Sigma},\operatorname{\mathbb{R}}^{2\times 2}_{sym})$ such that
@@
Comparing the proof of Proposition 3.1 in [4], the key ingredient and difference of the proof for our Proposition 3.1 is applying the elliptic estimate in Lemma 3.4 to control the new error terms with the phase function.
@@
Same as [4], we also divide the proof into three steps.
Step 1. Mollification. Take the lengthscale $\ell=\lambda^{-\tau}$, Set $\tilde{\rho}=\rho*\varphi_{\ell}$ and $\tilde{H}=H*\varphi_{\ell}$. Then by (3.7), (3.8) and Lemma A.1, we know,
@@
for $j\in\operatorname{\mathbb{N}}_{+}$. Let $\tilde{h}=\tilde{\rho}^{2}(\mathrm{Id}+\tilde{H})$ and $h=\rho^{2}(\mathrm{Id}+{H})$, then
@@
$+\tilde{H}\geq(1-\lambda^{-\alpha})\textrm{Id}>\tfrac{1}{2}\textrm{Id},$ Id ,
@@
(3.18) $\|\tilde{v}\|_{2+j}+\|\tilde{w}\|_{2+j}\leq C\ell^{-j}(\|v\|_{2}+\|w\|_{2})\leq C\delta^{\frac{1}{2}}\lambda^{1+\tau j},$
@@
where $\Gamma_{i},i=1,2$ are defined as follows.
@@
for $j=0,1,2.$ Here $\|\Gamma_{i}\|_{j}$ denotes $C^{j}$ norms of function $x\rightarrow\Gamma_{i}(a(x),\lambda^{\tau}\Phi_{1}(x)).$
@@
The proof of Lemma 3.2 is same as that in [4], but for the completeness, we put it in Appendix.
@@
(3.26) $\|v^{\lozenge}-v\|_{j}\leq C\delta^{1/2}\lambda^{\tau j-\tau},\,\,j=0,1,2.$
@@
Next we shall control $C^{j},j=0,1$ norms of the error matrix $\mathcal{E}.$ Taking derivatives gives
@@
$\mathcal{E}_{1}=$ $D(v^{\lozenge},w^{\lozenge})-(D(v,w)+a^{2}\nabla\Phi_{1}\otimes\nabla\Phi_{1})=\sum_{k=1}^{5}\mathcal{E}_{1k}$
@@
$\mathcal{E}_{11}:=$ $\lambda^{-\tau}(\partial_{t}\Gamma_{1}\partial_{s}\Gamma_{1}\mathrm{sym}(\nabla\Phi_{1}\otimes\nabla a)-\Gamma_{1}\nabla^{2}\tilde{v}+\partial_{s}\Gamma_{2}\textrm{sym}(\nabla\Phi_{1}\otimes\nabla a)),$
$\mathcal{E}_{12}:=$ ${\lambda}^{-\tau}\partial_{s}\Gamma_{1}\textrm{sym}((\nabla v-\nabla\tilde{v})\otimes\nabla a)+\partial_{t}\Gamma_{1}\textrm{sym}((\nabla v-\nabla\tilde{v})\otimes\nabla\Phi_{1}),$
@@
$\mathcal{E}_{15}:=$ $((v^{\lozenge}-v)\cot\Theta)\cdot\mathrm{Id}$
Estimates of the first three errors and their proof are same as those in [4].
@@
Assume all the assumptions of Proposition 3.1 are fulfilled, then for $j=0,1$, the following estimates hold.
@@
$\Delta(V^{\lozenge}-V)=F(v^{\lozenge})-F(v)&=\operatorname{div}\zeta+f,\mbox{ in }\Omega,\\ V^{\lozenge}-V&=0,\mbox{ on }\partial\Omega.\end{aligned}\right.$
Hence we require the follow elliptic estimates.
@@
where $\mathrm{diam}(\Omega)$ is the diameter of $\Omega$. Now (3.29) will follow by (3.30), (3.31) and the claim below:
@@
$\int_{\Omega}u_{1}^{2}\mathrm{d}x$ $=\int_{\Omega}u_{1}\Delta\phi\mathrm{d}x=\int_{\Omega}\phi\Delta u_{1}\mathrm{d}x=\int_{\Omega}\phi\operatorname{div}\zeta\mathrm{d}x$
$=-\int_{\Omega}\nabla\phi\cdot\zeta\mathrm{d}x\leq\|\nabla\phi\|_{L^{2}}\|\zeta\|_{L^{2}}.$

### v2

@@ Abstract.
We investigate the Dirichlet problem of the two-dimensional Lagrangian mean curvature equation in bounded domains. Infinitely many $C^{1,\alpha}(\alpha\in(0,\frac{1}{5}))$ very weak solutions are built through Nash-Kuiper construction. Moreover, we note there are infinitely many $C^{1,\alpha}$ very weak solutions that can not be improved to be $C^{2,\alpha}$.
2020 Mathematics Subject Classification:
35A02, 35D99, 35J25, 35J60, 35R25
@@ 1. Introduction
In the present paper, we study the Dirichlet problem for the two-dimensional Lagrangian mean curvature equation
@@
where $\Omega\subset\operatorname{\mathbb{R}}^{2}$ is a bounded domain, and the phase function $\Theta:\overline{\Omega}\to(-\pi,\pi)$ is a given $C^{2,\kappa}(\kappa\in(0,1))$ function. The equation
@@
(1.3) $\arctan\lambda_{1}+\arctan\lambda_{2}=\Theta$
@@
Generally, the $n$-dimensional Lagrangian mean curvature equation is
@@
where $\lambda_{i}$’s are eigenvalues of $\nabla^{2}v$. The mean curvature vector of the $n$-dimensional gradient graph $\Sigma$ is equal to $J\nabla_{g}\Theta$, where $g=\mathrm{Id}+(\nabla^{2}v)^{2}$ is the induced metric on $\Sigma$, and J(x,y)=(y,x)J(x,y)=(-y,x) is a rotation in $\operatorname{\mathbb{R}}^{n}\times\operatorname{\mathbb{R}}^{n}$. In particular, when $\Theta$ is a constant, (1.4) is called the special Lagrangian equation, which originates in the special Lagrangian geometry introduced by Harvey-Lawson [HL]. Note $\Sigma$ is a minimal surface in $\operatorname{\mathbb{R}}^{n}\times\operatorname{\mathbb{R}}^{n}$ if and only if $\Theta$ is a constant.
@@ 1.1. Background
In this paper, we consider not only classical solutions but also very weak solutions to (1.1). In fact, the very weak solution to the Monge-Ampère equation was first introduced by Lewicka-Pakzad [LePa] via very weak Hessian defined by Iwaniec [Iwan]. The two-dimensional very weak Hessian coincides with the Monge-Ampère measure for any convex solution ([DGG]). Moreover, Without boundary conditions, Lewicka-Pakzad [LePa], Székelyhidi and the first author [CSze], and Hirsch-Inauen and the first author [CHI] proved that there are infinitely many very weak solutions to the two dimensional Monge-Ampère equation of different regularity. After imposing boundary data, the first author in [Cao] studied very weak solutions to the Monge-Ampère equation. Recently, Li-Qiu [LQ] also constructed infinitely many very weak solutions of the Dirichlet problem for 2-Hessian equations. All the above approaches to constructing very weak solutions rely on the Nash-Kuiper construction, which was originated from isometric embeddings (the Nash-Kuiper theorem; see [Nash54, Kui55, CDS12, CS19-a, DIS, Sze]), and developed to be convex integration to prove the $h$-principle (see [Gro]). In recent decades, the technique of convex integration becomes a powerful approach to showing non-uniqueness of nonlinear partial differential equations.
@@ 1.2. Main result and strategies
Along the line of very weak solutions to the Monge-Ampère equation, we also introduce the concept of very weak solutions to the Dirichlet problem (1.1).
@@
$v\in C^{1}(\Omega)$ is a very weak solution to the Dirichlet problem (1.1) in $\Omega$, if $v=g$ on $\partial\Omega$, and it satisfies the following equation in the very weak sense in $\Omega$:
(1.5) $\cos\Theta\mbox{ curl curl }(v\cdot\mbox{Id})+\sin\Theta\left(-\frac{1}{2}\mbox{ curl curl }(\nabla v\otimes\nabla v)-1\right)=0;$
namely, for any $\phi\in C^{\infty}_{0}(\Omega)$,
@@
(1.6) $\quad+\int_{\Omega}v\Delta(\phi\cos\Theta)dx-\int_{\Omega}\phi\sin\Theta dx=0.$
The operator $\mathrm{curl}$ in two-dimension is defined by
$\mbox{ curl }\xi$ $=\partial_{2}\xi_{1}-\partial_{1}\xi_{2}\mbox{ for }\xi=(\xi_{1},\xi_{2})\in\mathbb{R}^{2},$
$\mbox{ curl }A$ $=(\partial_{2}A_{11}-\partial_{1}A_{12},\partial_{2}A_{21}-\partial_{1}A_{22})\mbox{ for }A=(A_{ij})_{2\times 2}\in\operatorname{\mathbb{R}}^{2\times 2}.$
In particular,  $\mbox{ curl curl }(v\cdot\mbox{Id})$ is understood as $\Delta v$ in the very weak sense.
@@
Let $\Omega\subset\operatorname{\mathbb{R}}^{2}$ be a bounded simply connected domain with a $C^{2,\kappa}$ boundary $\partial\Omega$ for any given $\kappa\in(0,1)$. Suppose $g\in C^{2,\kappa}(\partial\Omega)$ and the phase $\Theta:\overline{\Omega}\to(-\pi,\pi)$ is a $C^{2,\kappa}$ function. Then the following statements hold.
@@
    There exists some small positive number $c_{1}$ such that if $|\sin\Theta|\leq c_{1}$, then the Dirichlet problem (1.1) admits a unique $C^{2,\kappa}(\Omega)$ classical solution.
@@
    Given any constant $c_{2}\in(0,1]$, the Dirichlet problem (1.1) admits infinitely many $C^{1,\alpha}(\Omega)$ very weak solutions for any $\alpha\in(0,\frac{1}{5})$ whenever $|\sin\Theta|\geq c_{2}$.
@@
The assumption of simply connectedness is only needed for applying Lemma A.2 to the proof of Proposition 3.1.
Noting $c_{2}$ is arbitrary, as a quick observation, we have
@@
Our strategy to show Theorem 1.1 (1) is considering (1.2) as a perturbation of a Laplacian equation and then applying the implicit function theorem. As for Theorem 1.1 (2), with $|\sin\Theta|\geq c_{2}>0$, the problem is reduced to solve
@@
We will solve the above equation by the Nash-Kuiper construction. The main obstacle is how to treat terms with the phase function, compared with [Cao], which will be regarded as a Laplacian of some function, i.e.,
$\cot\Theta\Delta v=\Delta(v\cot\Theta-V),\text{ in the very weak sense,}$
@@
$\mathrm{curl}\,\mathrm{curl}\,\big(\mbox{ sym }\nabla w+\frac{1}{2}\nabla v\otimes\nabla v-(v\cot\Theta)\cdot\mathrm{Id}+V\cdot\mathrm{Id}\big)=-1,$
in the very weak sense, and then $V\cdot\textrm{Id}$ becomes an error term of a lower order. With the elliptic estimate in Lemma 3.4, we are able to control it and arrive at our conclusion.
We remark that the method proving Theorem 1.1 (2) has more general applications.
@@
where $a=\cot\tau$, and $b=\sqrt{|\cot^{2}\tau-1|}$. With slight modifications, we can establish the result like Theorem 1.1 (2) for the following Dirichlet problem
@@
where $\lambda_{1}$, $\lambda_{2}$ are eigenvalues of $\nabla^{2}v$. We note the equation $F_{\tau}(\lambda)=\Theta$ can be formulated as the equation of the following form
@@
Equations in (1.7) are originated from the study of the gradient graph $\Sigma$. By Warren [Wa], if $\Sigma$ is a minimal Lagrangian graph in $(\operatorname{\mathbb{R}}^{2}\times\operatorname{\mathbb{R}}^{2},g_{\tau})$ with
$g_{\tau}=\sin\tau\big(\sum_{i=1}^{2}dx_{i}\otimes dx_{i}+\sum_{j=1}^{2}dy_{j}\otimes dy_{j}\big)+\cos\tau\big(\sum_{i=1}^{2}dx_{i}\otimes dy_{i}+\sum_{j=1}^{2}dy_{j}\otimes dx_{j}\big),$
@@
$F_{\tau}(\lambda(\nabla^{2}v))=\text{constant}.$
We refer readers to the results about gradient graphs in general dimensions in [Wa].
@@ 1.3. Organization and notations
$h]_{m+\alpha}=\max_{|\beta|=m}\sup_{x\neq y}\frac{|\partial^{\beta}h(x)-\partial^{\beta}h(y)|}{|x-y|^{\alpha}}.$
Here $m$ is a nonnegative integer, $\alpha\in(0,1]$ a constant, and $\beta$ be a multi-index. We denote by $B_{r}$ the ball in $\operatorname{\mathbb{R}}^{2}$ with radius $r$ centering at 0. Let $\varphi\in C^{\infty}_{0}(B_{1})$ be a standard two-dimensional mollifier. We define $\varphi_{l}(x)=l^{-2}\varphi(l^{-1}x)$, and $l>0$ is called the length-scale. We also define
$\tilde{h}(x)=h*\varphi_{l}(x)=l^{-2}\int_{\operatorname{\mathbb{R}}^{2}}h(x-y)\varphi(\frac{y}{l})dy.$
Some properties of mollification can be found in Lemma A.1.
Besides, for $\xi=(\xi_{1},\xi_{2})\in\mathbb{R}^{2}$ and $\eta=(\eta_{1},\eta_{2})\in\mathbb{R}^{2}$, we write $\xi\otimes\eta$ as a $2\times 2$-matrix with entries $(\xi\otimes\eta)_{ij}=\xi_{i}\eta_{j}$. By a direct calculation, for a $C^{3}$-function $w$, we have
@@
Acknowledgements
Wentao Cao’s research was supported in part by National Natural Science Foundation of China (Grant No. 12471224). Zhehui Wang’s research was supported in part by Young Scientists Fund of National Natural Science Foundation of China (Grant No. 12401247), and Guangdong Basic and Applied Basic Research Foundation (Grant No. 2023A1515110910).
@@ Acknowledgements
The proof of Theorem 1.1 (1) is based on the application of the implicit function theorem.
@@
Claim: There are a unique $\psi\in\mathcal{X}$ and a small enough constant $\mu\in(0,1)$ such that $G(\psi,\Psi)=0$ for any $\Psi$ satisfying $\|\Psi\|_{\kappa}\leq\mu$.
@@
for any $0<\sigma\leq r$ and $f\in C^{1,\kappa}(B_{r})$. Hence, if $|\sin\Theta|\leq c_{1}$ with the small constant $c_{1}$ to be fixed, then we can get
@@
Then, $v:=\psi+\widetilde{g}$ uniquely satisfies
$\left\{\begin{aligned} &\Delta v+\tan\Theta(\det\nabla^{2}v-1)=0\mbox{ in }\Omega,\\ &v=g\mbox{ on }\partial\Omega.\end{aligned}\right$.
@@
3. The Stage Proposition
The rest of the paper is devoted to showing Theorem 1.1 (2). Similar to the treatment in [Cao, CSze, LePa] for the Monge-Ampère equation, we split (1.2) (after dividing $\sin\Theta$) into a differential system
(3.1) $-\mbox{curl curl }A=1,$
@@ 3. The Stage Proposition
$\left\{\begin{aligned} \Delta V&=F(v),\mbox{ in }\Omega,\\ V&=0,\mbox{ on }\partial\Omega,\end{aligned}\right$.
@@
In fact, by a direct calculation, we note that $v$ satisfying (3.1)-(3.2) is a solution of the equation (1.5) in the very weak sense.
To solve (3.2), our strategy is applying the technique of Nash-Kuiper construction to construct a sequence of subsolutions satisfying the boundary condition, whose limit is our desired solution. Set
@@
Note that the definition is a slight modification of that in [Cao]. To construct new adapted subsolutions, we need to add the deficit AD(v,w)A-D(v,w) to the given adapted subsolution $v.$ Since the deficit is of the form $\rho^{2}(\mathrm{Id}+H)$, we require the following stage proposition (in the terminology of Nash [Nash54]) for iteration, which is parallel to Proposition 3.1 in [Cao] and whose proof is different due to the new error terms. The new error terms are then controlled through Lemma 3.4.
@@
Suppose $\Sigma\subset\operatorname{\mathbb{R}}^{2}$ is a simply connected open bounded set with a $C^{2}$ boundary. Let $v\in C^{2}(\overline{\Sigma})$ and $w\in C^{2}(\Sigma,\operatorname{\mathbb{R}}^{2})$ such that
@@
and let $\rho\in C^{1}(\overline{\Sigma})$, $H\in C^{1}(\overline{\Sigma},\operatorname{\mathbb{R}}^{2\times 2}_{sym})$ be such that
@@
for some $\gamma>1$, $0<\delta,\,\alpha<1$ and $\lambda>1$ with $2\leq\lambda^{\alpha}$ and $\delta^{1/2}\lambda>1$. Then for any $\tau>1$, there exist functions $v^{\star}\in C^{2}(\overline{\Sigma})$, $w^{\star}\in C^{2}(\overline{\Sigma},\operatorname{\mathbb{R}}^{2})$ and a symmetric matrix $\operatorname{\mathcal{E}}\in C^{1}(\overline{\Sigma},\operatorname{\mathbb{R}}^{2\times 2}_{sym})$ such that
@@
Our proof for the proposition above differs from that for Proposition 3.1 in [Cao] due to the presence of the phase function $\Theta$. Indeed, the key ingredients of our proof consist of applying the elliptic estimates in Lemma 3.4 to control the error terms arising from $\Theta$.
@@
We divide the proof into three steps.
Step 1. Mollification. Take the lengthscale $\ell=\lambda^{-\tau}$. Set $\tilde{\rho}=\rho*\varphi_{\ell}$ and $\tilde{H}=H*\varphi_{\ell}$. Then by (3.7), (3.8) and Lemma A.1, we know,
@@
[…]
@@
$\textrm{Id}+\tilde{H}\geq(1-\lambda^{-\alpha})\textrm{Id}>\tfrac{1}{2}\textrm{Id},$
@@
(3.18) $\|\tilde{v}\|_{2+j}+\|\tilde{w}\|_{2+j}\leq C\ell^{-j}(\|v\|_{2}+\|w\|_{2})\leq C\delta^{\frac{1}{2}}\lambda^{1+\tau j}.$
@@
where $\Gamma_{i},i=1,2$ are defined as follows:
@@
for $j=0,1,2.$ Here $\|\Gamma_{i}\|_{j}$ denotes the $C^{j}$ norms of the function $x\rightarrow\Gamma_{i}(a(x),\lambda^{\tau}\Phi_{1}(x)).$
@@
The proof of Lemma 3.2 is the same as that in [Cao], but for the completeness, we put it in Appendix.
@@
(3.26) $\|v^{\lozenge}-v\|_{j}\leq C\delta^{1/2}\lambda^{\tau j-\tau},\,\,j=0,1,2,$
@@
Next we shall control the $C^{j}$($j=0,1$) norms of the error matrix $\mathcal{E}.$ Taking derivatives gives
@@
$\mathcal{E}_{1}=$ $D(v^{\lozenge},w^{\lozenge})-\left[D(v,w)+a^{2}\nabla\Phi_{1}\otimes\nabla\Phi_{1}\right]=\sum_{k=1}^{5}\mathcal{E}_{1k}$
@@
$\mathcal{E}_{11}:=$ $\lambda^{-\tau}\left[\partial_{t}\Gamma_{1}\partial_{s}\Gamma_{1}\mathrm{sym}(\nabla\Phi_{1}\otimes\nabla a)-\Gamma_{1}\nabla^{2}\tilde{v}+\partial_{s}\Gamma_{2}\textrm{sym}(\nabla\Phi_{1}\otimes\nabla a)\right],$
$\mathcal{E}_{12}:=$ ${\lambda}^{-\tau}\partial_{s}\Gamma_{1}\textrm{sym}\left[(\nabla v-\nabla\tilde{v})\otimes\nabla a\right]+\partial_{t}\Gamma_{1}\textrm{sym}\left[(\nabla v-\nabla\tilde{v})\otimes\nabla\Phi_{1}\right],$
@@
$\mathcal{E}_{15}:=$ $\left[(v^{\lozenge}-v)\cot\Theta\right]\cdot\mathrm{Id}$
Estimates of the first three errors and their proofs are the same as those in [Cao].
@@
Assume all the assumptions of Proposition 3.1 are fulfilled, then for $j=0,1$, the following estimates hold:
@@
$\left\{\begin{aligned} \Delta(V^{\lozenge}-V)=F(v^{\lozenge})-F(v)&=\operatorname{div}\zeta+f,\mbox{ in }\Omega,\\ V^{\lozenge}-V&=0,\mbox{ on }\partial\Omega.\end{aligned}\right$.
Hence we require the following elliptic estimates.
@@
where $\mathrm{diam}(\Omega)$ is the diameter of $\Omega$. Now (3.29) will follow from (3.30), (3.31) and the claim below:
@@
$\int_{\Omega}u_{1}^{2}dx$ $=\int_{\Omega}u_{1}\Delta\phi dx=\int_{\Omega}\phi\Delta u_{1}dx=\int_{\Omega}\phi\operatorname{div}\zeta dx$
$=-\int_{\Omega}\nabla\phi\cdot\zeta dx\leq\|\nabla\phi\|_{L^{2}}\|\zeta\|_{L^{2}}.$